6a.) TTM: 10 years. YTM: 5%. C: 5%. PV:? 6b.) TTM: 10 years. YTM: 7%. C: 5%. PV:? 6c.) TTM: 10 years. YTM: 3%. C: 5%. PV:? 7a.) TTM: 10 years. YTM: 11%. C: 5%. PV:? 7b.) TTM: 10 years. YTM: 13% C: 5% PV:? 7c.) TTM: 10 years. YTM: 9% C: 5%. PV: ? Questions for Problems 6-7 a.) What do problems 6 and 7 tell us about the way the initial yield to maturity affects the sensitivity of a bonds price to a change in the level of interest rates? b.) Why is this the result?
Answer #1

Question 6:

a. PV is calculated using =PV(rate,nper,pmt,fv) in excel =PV(0.05,10,50,1000) = $1,000

b. PV =PV(0.07,10,50,1000) = $859.53

c. PV =PV(0.03,10,50,1000) = $1,170.60

This shows us that there is an inverse relationship between the yield to maturity and the bond price. As we see in #6 above, as the YTM increases above the coupon rate, the bond trades at a discount and as the YTM decreases below the coupon rate, the bond trades at a premium

Note: We have answered one question (#6) with all sub-parts (a,b and c). Please post #7 separately for experts to answer since only one full question can be answered at a time.